# Near-linear Time Dispersion of Mobile Agents

### v2

@@ 4.2 Details
Thereafter, $a$ invokes $\mathtt{Probe}()$ at the end of slot 2. This subroutine runs in slots 4–8. While $\mathtt{Probe}()$ in $\mathbf{RootedDispersion}$ returns the control to the main function after completing the probing, i.e., determining whether or not an unsettled neighbor exists, $\mathtt{Probe}()$ in $\mathbf{GeneralDispersion}$ returns the control each time slot 8 ends even if it does not complete the probing. Consider that there are $x-1$ accompanying zombies when a leader $a$ begins the probing. First, a leader $a$ and the $x-1$ accompanying zombies join the probing. Each of them, say $b$, moves from a node $w$ to one of its neighbors $u\in N(w)$ in slot 6 (line 47) and goes back to $w$ in slot 7 (line 54). If $b$ finds a settler in the same group at $u$, it sets $\psi(u).\mathtt{help}$ to $b.\mathtt{pin}$ (line 51). As long as $s.\mathtt{help}\neq\bot$, a settler $s$ at a node $v$ goes to a neighbor $N(v,s.\mathtt{help})$ in slot 3 (line 60, Algorithm 4). Hence, in the next slot 3, that settler $\psi(u)$ goes to $w$. If there are $2x$ agents at $w$ excluding $\psi(w)$, those $2x$ agents perform the same process in the next slots 6 and 7, that is, they go to unprobed neighbors, update the $\mathtt{help}$ of settlers in the same group (if exists), and go back to $w$. In slot 8, $a$ sends the helping settlers back to their home. The number of agents joining the probing at $w$, i.e.$|A(w)\setminus\{\psi(w)\}|$, doubles at each iteration of this process until they find a node without a settler in the same group or finish probing all neighbors in $N(w)$. Thus, like $\mathbf{RootedDispersion}$, the probing finishes in $O(\log\tau)$ time steps. At this time, $\psi(w).\mathtt{next}=\bot$ holds if all neighbors in $N(w)$ are settled by settlers in the same group. Otherwise, $N(w,\psi(w).\mathtt{next})$ is unsettled or settled by a settler in another group. Then, in the next slot 5, $a$ resets the $\mathtt{help}$ of all settlers at $w$ to $\bot$ except for $\psi(w)$, lets them go back to their homes, and sets $\psi(w).\mathtt{done}$ to $\mathbf{true}$, indicating that the probing is done (lines 38–40). The probing process described above may be prevented by a stronger leader $b$ when $b$ visits a node $v\in N(w)$ such that $\psi(v)$ belongs to $a$’s group and $\psi(v).\mathtt{help}\neq\bot$. Then, $b$ incorporates $\psi(v)$ into $b$’s group, and set $\psi(v).\mathtt{help}$ to $\bot$ (line 19), so $\psi(v)$ never goes to $w$ to help $a$’s probing. However, this event actually speeds up $a$’s probing: $a$ identifies this event when noticing that $\psi(v)$ does not arrive at $w$ in the next slot 4. As a result, $a$ can set $\psi(w).\mathtt{next}$ to $p$ where $N(w,p)=v$ (lines 32–34).
@@
If a leader $a$ at $w$ observes $\psi(w).\mathtt{done}=\mathbf{true}$, it makes a forward or backward move in slot 11 (lines 22–29). Each time $a$ makes a forward or backward move to a node $u$, it remembers $a.\mathtt{pin}$ in $a.\mathtt{parent}$ after the move (line 28). This port number will be stored on the variable $\psi(u).\mathtt{parent}$ when $a$ settles a zombie on $u$ or $a$ incorporates $\psi(u)$ from the territory of another group. Note that this event occurs only when the last move is forward. Thus, like $\mathbf{RootedDispersion}$, $a$ constructs a DFS tree in its territory. It is inevitable to use a variable $a.\mathtt{parent}$ tentatively since $a.\mathtt{pin}$ is updated every step by definition of a special variable $\mathtt{pin}$ and $a$ may become a waiting leader after moving to $u$. Unlike $\mathbf{RootedDispersion}$, $a$ records the most recently used port to move in $\psi(\nu(a)).\mathtt{next}$ even when it makes a backward move (lines 24–25). This allows a zombie to chase a leader.
@@
Neither a leader nor a settler decreases its level in $\mathbf{GeneralDispersion}$. When a zombie $z$ does not accompany a leader, it chases a leader through port $\psi(\nu(z)).\mathtt{next}$. This port $\psi(\nu(z)).\mathtt{next}$ is updated only if a leader makes a forward or backward move from $\nu(z)$, and the leader updates the level of $\psi(N(\nu(z),\psi(\nu(z)).\mathtt{next}))$ if it is smaller than its level. Thus, a zombie never decreases its location level by chasing a leader. When a zombie $z$ accompanies a leader, the leader copies its level to $\psi(\nu(z)).\mathtt{level}$ in slot 2, which is copied to $z.\mathtt{level}_{L}$ in slot 8. The leader that $z$ accompanies may change but does not change to a weaker leader. Thus, a zombie never decreases its location level when accompanying a leader. ∎
@@
Since an agent in $\mathbf{GeneralDispersion}$ manages only a constant number of variables, each with $O(\log(k+\delta))$ bits, Lemmas 3 and 6 yield the following theorem.
Theorem 2.
In the general setting, there exists an algorithm that solves the dispersion problem within $O(k(\log\tau))(\log k))=O(k\log^{2}k)$ time using $O(\log(k+\delta))$ bits of space per agent.
5 For Further Improvement in Time Complexities
1 $\mathtt{Settle}(b,s,\bot)$, where $s$ is the starting node and $b$ is the agent with the smallest ID in $A(s)$ $\mathtt{Probe}()$ while $a_{\mathrm{max}}.\mathtt{settled}=\bot$ do
2       Let $w=\nu(a_{\mathrm{max}})$ if $U_{0}(w)\neq\bot$ then
             Chose any neighbor $u\in U_{0}(w)$   // $u$ is an unsetteled node here
3             All explorers in $A(w)$ go to $u$ $\mathtt{Settle}(b,u,a_{\mathrm{max}}.\mathtt{pin})$, where $b^{\prime}$ is the agent with the smallest ID in $A(u)$ $\mathtt{Probe}()$
@@ 2       Let $w=\nu(a_{\mathrm{max}})$ if $U_{0}(w)\neq\bot$ then
6
7function $\mathtt{Settle}(b,v,p)$:
8       $b.\mathtt{settled}\leftarrow\top$ $b.\mathtt{parent}\leftarrow p$ for $i\in[0,\delta_{v}]$ do
@@
10
11function $\mathtt{Probe}()$:
       Let $w=\nu(a_{\mathrm{max}})$ Let $\ell=|A(w)\setminus\{\psi(w)\}|$   // $\ell$ is the number of explorers
@@
             Let $a_{1},a_{2},\dots,a_{x}$ be the agents in $A(w)\setminus\{\psi(w)\}$  // $x\leq\ell$
13             Let $\delta^{\prime}=\min(x,|U_{\bot}(w)|)$ Choose $\delta^{\prime}$ neighbors $u_{1},u_{2},\dots,u_{\delta^{\prime}}$ from $U_{\bot}$ arbitrarily. for each $a_{i}\in\{a_{1},a_{2},\dots,a_{\delta^{\prime}}\}$ in parallel do
14                   $a_{i}$ moves to $u_{i}$ if $a_{i}$ finds $\psi(u_{i})$ at $u_{i}$ then
@@
16                  $a_{i}$ goes back to $w$ along with $\psi(u_{i})$ (if it exists) Let $p_{i}$ be a port number such that $N(w,p_{i})=u_{i}$ $\psi(w).\mathtt{checked}[p_{i}]\leftarrow\begin{cases}1&\text{if }a_{i}\text{ found }\psi(u_{i})\text{ at }u_{i}\\ 0&\text{otherwise},\end{cases}$
17
18      Let all settlers except for $\psi(w)$ go back to their homes if $|U_{0}(w)|\geq\ell$ then
@@
20
Algorithm 6 $\mathbf{RootedOptimal}$
@@
In the previous sections, we introduced nearly time-optimal dispersion algorithms: an $O(k\log\tau)$-time algorithm for the rooted setting and an $O(k\log^{2}k)$-time algorithm for the general setting. This raises a crucial question: is it possible to develop a truly time-optimal algorithm, specifically an $O(k)$-time algorithm, even if it requires much more space? In this section, we affirmatively answer this question for the rooted setting. We present an $O(k)$-time algorithm that utilizes $O(\delta)$ bits of space per agent. However, the feasibility of an $O(k)$-time algorithm in the general setting remains open.
@@
We refer to the new algorithm as $\mathbf{RootedOptimal}$ in this section. In $\mathbf{RootedDispersion}$, with $O(\log\delta)$ bits of space, each settler $\psi(w)$ cannot memorize the exact set of settled neighbors of $w$. Instead, it only remembers that the first $i$ neighbors of $w$ are settled for some $i$. In contrast, $\mathbf{RootedOptimal}$ allows each settler $\psi(w)$ to remember all settled neighbors of $w$ using $O(\delta)$ bits, which significantly helps to eliminate an $O(\log\tau)$ factor from the time complexity. However, somewhat surprisingly, both the design of the new algorithm and the analysis of its execution time are non-trivial.
@@
Below, we outline the modifications made to $\mathbf{RootedDispersion}$ to obtain $\mathbf{RootedOptimal}$. The pseudocode for $\mathbf{RootedOptimal}$ is provided in Algorithm 6. However, we anticipate that most readers will grasp the behavior of $\mathbf{RootedOptimal}$ simply by reviewing the following key differences.
@@
  • In $\mathbf{RootedOptimal}$, each settler $\psi(w)$ maintains an array variable $\psi(w).\mathtt{checked}$ of size $\delta_{w}$. Each element of $\psi(w).\mathtt{checked}[i]$ takes a value from the set $\{0,1,\bot\}$. The assignment $\psi(w).\mathtt{checked}[i]=0$ (respectively, $1$) indicates that the neighbor $N(w,i)$ is unsettled (respectively, settled). The value $\bot$ is utilized exclusively during the probing process, meaning that the neighbor $N(w,i)$ has yet to be checked for its settled status. For any $c\in\{0,1,\bot\}$, we define $U_{c}=\{N(u,i)\mid i\in[0,\delta_{w}-1],\psi(w).\mathtt{checked}[i]=c\}$.
@@
  • Consider that the unique leader $a_{\mathrm{max}}$ invokes the probing process $\mathtt{Probe}()$ at a node $w$. (Remember that $a_{\mathrm{max}}$ is the unique leader that has the maximum identifier at the beginning of the execution.) In $\mathbf{RootedDispersion}$, the probing process terminates as soon as any agent finds an unsettled neighbor. However, in $\mathbf{RootedOptimal}$, the process only ends when all of $w$’s neighbors are probed or when at least $\ell$ unsettled neighbors are found, where $\ell$ is the number of explorers. In the former case, the probing process is now complete: $U_{\bot}(w)$ is empty, and $U_{0}(w)$ equals the set of unsettled neighbors of $w$. In the latter case, the explorers go to distinct $\ell$ unsettled nodes and settle there, thereby achieving dispersion.
@@
  • Consider an agent $a$ making a round trip $w\to u\to w$ during $\mathtt{Probe}()$, where $u=N(w,p)$ for some $p\in[0,\delta_{w}-1]$. If $a$ does not encounter a settler at $u$, it simply sets $\psi(w).\mathtt{checked}[i]$ to 00. On the other hand, if a settler is found at $u$, $a$ sets $\psi(w).\mathtt{checked}[i]$ to $1$ and additionally sets $\psi(u).\mathtt{checked}[q]$ to $1$, where $q\in[0,\delta_{u}-1]$ is the port number such that $w=N(u,q)$. This modification ensures that $U_{0}(u)$ remains equal to the set of unsettled neighbors of $u$ when $\mathtt{Probe}$ finishes.
@@
  • In $\mathbf{RootedDispersion}$, the leader $a_{\mathrm{max}}$ invokes $\mathtt{Probe}()$ after each forward or backward move. However, in $\mathbf{RootedOptimal}$, $a_{\mathrm{max}}$ only invokes $\mathtt{Probe}()$ after making a forward move. This change does not compromise the correctness of $\mathbf{RootedOptimal}$ because when $a_{\mathrm{max}}$ makes a backward move to a node $w$, $\psi(w)$ accurately remembers its unsettled neighbors due to the modification mentioned earlier.
@@
One might think that the probing process $\mathtt{Probe}()$ in $\mathbf{RootedOptimal}$ could take longer time than in $\mathbf{RootedDispersion}$, as it only finishes after all neighbors of the current location have been probed or after finding $\ell$ unsettled neighbors, where $\ell$ is the number of explorers. Particularly, there seems to be a concern that during the probing at a node $w$, the number of agents, excluding $\psi(w)$, may not always double: this event occurs when some agents discover an unsettled neighbor. Despite that, we deny this conjecture at least asymptotically, that is, we have the following lemma.
@@
Assume that $a_{\mathrm{max}}$ invokes $\mathtt{Probe}()$ at node $w$ during the execution of $\mathbf{RootedOptimal}$, and exactly $\ell$ explorers including $a_{\mathrm{max}}$ exists at the time. Then, $\mathtt{Probe}()$ finishes within $O(1)+\max(0,2\lceil\log\tau-\log\ell\rceil)$ time.
@@
If $\ell\geq\tau$, we have $\ell\geq\min(\delta,k)=\delta$ because $\ell<k$. Then, the lemma trivially holds: $\mathtt{Probe}()$ finishes in a constant time. Thus, we consider the case $\ell<\tau$. Let $t$ be the time step at which $a_{\mathrm{max}}$ invokes $\mathtt{Probe}()$ at a node $w$, and let $z=\lceil\log\tau-\log\ell\rceil+1$. It suffices to show that $U_{\bot}(w,t^{\prime})=\emptyset$ or $U_{0}(w,t^{\prime})\geq\ell$ holds for some $t^{\prime}\in[t,t+2z]$. Assume for contradiction that this does not hold. For any $r\in[0,z]$, we define
$f(r)=X_{r}\cdot 2^{r-1}+Y_{r},$
where $X_{r}=|U_{0}(w,t+2r)|$ and $Y_{r}=|A(w,t+2r)\setminus\{\psi(w)\}|$. By definition, $f(0)=0\cdot 2^{0-1}+\ell=\ell$. Under the above assumption, for any $r=0,1,\dots,z-1$, the agents in $A(w,t+2r)$ move to distinct neighbors in $U_{\bot}(w,t+2r)$ in time step $t+2r$, and bring back all settlers they find, at most one for each neighbor, in time step $t+2r+1$. Let $\alpha$ be the number of those settlers i.e.$\alpha=Y_{r+1}-Y_{r}$. Note that $X_{r+1}-X_{r}=Y_{r}-\alpha$ hods here. Then, irrespective of $\alpha$, we obtain
$f(r+1)$ $=X_{r+1}\cdot 2^{r}+Y_{r+1}$
$=(X_{r}+Y_{r}-\alpha)2^{r}+Y_{r}+\alpha$
$=X_{r}\cdot 2^{r}+(2^{r}-1)(Y_{r}-\alpha)+2\cdot Y_{r}$
$\geq 2(x_{r}\cdot 2^{r-1}+Y_{r})$
$=2f(r),$
where we use $2^{r}\geq 1$ and $Y_{r}\geq\alpha$ in the above inequality. Therefore, we have $f(z)\geq\ell\cdot 2^{z}$, while $f(z)=X_{z}\cdot 2^{z-1}+Y_{z}\leq(\ell-1)2^{z-1}+Y_{z}$ by the assumption that $X_{z}=|U_{0}(w,t+2z)|<\ell$. Thus, we have $|A(w,t+2z)\setminus\{\psi(w)\}|=Y_{z}\geq\ell\cdot 2^{z}-(\ell-1)2^{z-1}=(\ell+1)2^{z-1}\geq(\ell+1)\tau/\ell>\tau$. Since $\tau=\min(\delta,k)$, we have $\tau=\delta$ or $\tau=k$. In the former case, $Y_{k}>\delta$ agents at $w$ are enough to visit all neighbors in $U_{\bot}(w,t+2z)$ in time step $t+2z$, thus $U_{\bot}(w,t+2z+2)=\emptyset$ holds, a contradiction. In the latter case, there are $|A(w,t+2z)|\geq k+2$ agents in $w$ at time step $t+2z$, a contradiction. Therefore, $U_{\bot}(w)=\emptyset$ or $U_{0}(w)\geq\ell$ holds at time step $t^{\prime}\leq t+2z+2$. ∎
@@
In the rooted setting, algorithm $\mathbf{RootedOptimal}$ solves the dispersion problem within $O(k)$ time using $O(\delta)$ bits of space per agent.

### v3

@@ 4.3 Detail Implementation of General Dispersion
Thereafter, $a$ invokes $\mathtt{Probe}()$ at the end of slot 2. This subroutine runs in slots 4–8. While $\mathtt{Probe}()$ in $\mathbf{RootedDisp}$ returns the control to the main function after completing the probing, i.e., determining whether or not an unsettled neighbor exists, $\mathtt{Probe}()$ in $\mathbf{GeneralDisp}$ returns the control each time slot 8 ends even if it does not complete the probing. Consider that there are $x-1$ accompanying zombies when a leader $a$ begins the probing. First, a leader $a$ and the $x-1$ accompanying zombies join the probing. Each of them, say $b$, moves from a node $w$ to one of its neighbors $u\in N(w)$ in slot 6 (line 47) and goes back to $w$ in slot 7 (line 54). If $b$ finds a settler in the same group at $u$, it sets $\psi(u).\mathtt{help}$ to $b.\mathtt{pin}$ (line 51). As long as $s.\mathtt{help}\neq\bot$, a settler $s$ at a node $v$ goes to a neighbor $N(v,s.\mathtt{help})$ in slot 3 (line 60, Algorithm 4). Hence, in the next slot 3, that settler $\psi(u)$ goes to $w$. If there are $2x$ agents at $w$ excluding $\psi(w)$, those $2x$ agents perform the same process in the next slots 6 and 7, that is, they go to unprobed neighbors, update the $\mathtt{help}$ of settlers in the same group (if exists), and go back to $w$. In slot 8, $a$ sends the helping settlers back to their home. The number of agents joining the probing at $w$, i.e.$|A(w)\setminus\{\psi(w)\}|$, doubles at each iteration of this process until they find a node without a settler in the same group or finish probing all neighbors in $N(w)$. Thus, like $\mathbf{RootedDisp}$, the probing finishes in $O(\log\tau)$ time steps. At this time, $\psi(w).\mathtt{next}=\bot$ holds if all neighbors in $N(w)$ are settled by settlers in the same group. Otherwise, $N(w,\psi(w).\mathtt{next})$ is unsettled or settled by a settler in another group. Then, in the next slot 5, $a$ resets the $\mathtt{help}$ of all settlers at $w$ to $\bot$ except for $\psi(w)$, lets them go back to their homes, and sets $\psi(w).\mathtt{done}$ to $\mathbf{true}$, indicating that the probing is done (lines 38–40). The probing process described above may be prevented by a stronger leader $b$ when $b$ visits a node $v\in N(w)$ such that $\psi(v)$ belongs to $a$’s group and $\psi(v).\mathtt{help}\neq\bot$. Then, $b$ incorporates $\psi(v)$ into $b$’s group, and set $\psi(v).\mathtt{help}$ to $\bot$ (line 19), so $\psi(v)$ never goes to $w$ to help $a$’s probing. However, this event actually speeds up $a$’s probing: $a$ identifies this event when noticing that $\psi(v)$ does not arrive at $w$ in the next slot 4. As a result, $a$ can set $\psi(w).\mathtt{next}$ to $p$ where $N(w,p)=v$ (lines 32–34).
@@
If a leader $a$ at $w$ observes $\psi(w).\mathtt{done}=\mathbf{true}$, it makes a forward or backward move in slot 11 (lines 22–29). Each time $a$ makes a forward or backward move to a node $u$, it remembers $a.\mathtt{pin}$ in $a.\mathtt{parent}$ after the move (line 28). This port number will be stored on the variable $\psi(u).\mathtt{parent}$ when $a$ settles a zombie on $u$ or $a$ incorporates $\psi(u)$ from the territory of another group. Note that this event occurs only when the last move is forward. Thus, like $\mathbf{RootedDisp}$, $a$ constructs a DFS tree in its territory. It is inevitable to use a variable $a.\mathtt{parent}$ tentatively since $a.\mathtt{pin}$ is updated every step by definition of a special variable $\mathtt{pin}$ and $a$ may become a waiting leader after moving to $u$. Unlike $\mathbf{RootedDisp}$, $a$ records the most recently used port to move in $\psi(\nu(a)).\mathtt{next}$ even when it makes a backward move (lines 24–25). This allows a zombie to chase a leader.
@@
Neither a leader nor a settler decreases its level in $\mathbf{GeneralDisp}$. When a zombie $z$ does not accompany a leader, it chases a leader through port $\psi(\nu(z)).\mathtt{next}$. This port $\psi(\nu(z)).\mathtt{next}$ is updated only if a leader makes a forward or backward move from $\nu(z)$, and the leader updates the level of $\psi(N(\nu(z),\psi(\nu(z)).\mathtt{next}))$ if it is smaller than its level. Thus, a zombie never decreases its location level by chasing a leader. When a zombie $z$ accompanies a leader, the leader copies its level to $\psi(\nu(z)).\mathtt{level}$ in slot 2, which is copied to $z.\mathtt{level}_{L}$ in slot 8. The leader that $z$ accompanies may change but does not change to a weaker leader. Thus, a zombie never decreases its location level when accompanying a leader. ∎
4.4 Proofs of Theorem 2
@@
Since an agent in $\mathbf{GeneralDisp}$ manages only a constant number of variables, each with $O(\log(k+\Delta))$ bits, Lemmas 3 and 6 yield Theorem 2.
1 $\mathtt{Settle}(b,s,\bot)$, where $s$ is the starting node and $b$ is the agent with the smallest ID in $A(s)$
2 $\mathtt{Probe}()$
3 while $a_{\mathrm{max}}.\mathtt{settled}=\bot$ do
4       Let $w=\nu(a_{\mathrm{max}})$
5       if $U_{0}(w)\neq\bot$ then
             Chose any neighbor $u\in U_{0}(w)$
              // $u$ is an unsetteled node here
6             All explorers in $A(w)$ go to $u$
7             $\mathtt{Settle}(b,u,a_{\mathrm{max}}.\mathtt{pin})$, where $b^{\prime}$ is the agent with the smallest ID in $A(u)$
8             $\mathtt{Probe}()$
@@
11
12
13
14function $\mathtt{Settle}(b,v,p)$:
15       $b.\mathtt{settled}\leftarrow\top$
16       $b.\mathtt{parent}\leftarrow p$
17       for $i\in[0,\delta_{v}]$ do
@@
19
20
21function $\mathtt{Probe}()$:
22       Let $w=\nu(a_{\mathrm{max}})$
       Let $\ell=|A(w)\setminus\{\psi(w)\}|$
        // $\ell$ is the number of explorers
@@ 22       Let $w=\nu(a_{\mathrm{max}})$
             Let $a_{1},a_{2},\dots,a_{x}$ be the agents in $A(w)\setminus\{\psi(w)\}$
              // $x\leq\ell$
24             Let $\Delta^{\prime}=\min(x,|U_{\bot}(w)|)$
25             Choose $\Delta^{\prime}$ neighbors $u_{1},u_{2},\dots,u_{\Delta^{\prime}}$ from $U_{\bot}$ arbitrarily.
26             for each $a_{i}\in\{a_{1},a_{2},\dots,a_{\Delta^{\prime}}\}$ in parallel do
27                   $a_{i}$ moves to $u_{i}$
28                   if $a_{i}$ finds $\psi(u_{i})$ at $u_{i}$ then
@@ 27                   $a_{i}$ moves to $u_{i}$
30
31                  $a_{i}$ goes back to $w$ along with $\psi(u_{i})$ (if it exists)
32                   Let $p_{i}$ be a port number such that $N(w,p_{i})=u_{i}$
33                   $\psi(w).\mathtt{checked}[p_{i}]\leftarrow\begin{cases}1&\text{if }a_{i}\text{ found }\psi(u_{i})\text{ at }u_{i}\\ 0&\text{otherwise},\end{cases}$
34
35
36      Let all settlers except for $\psi(w)$ go back to their homes
37       if $|U_{0}(w)|\geq\ell$ then
@@ 36      Let all settlers except for $\psi(w)$ go back to their homes
39
Algorithm 6 $\mathbf{RootedOpt}$
5 For Further Improvement in Time Complexities
@@
In the previous sections, we introduced nearly time-optimal dispersion algorithms: an $O(k\log\tau)$-time algorithm for the rooted setting and an $O(k\log^{2}k)$-time algorithm for the general setting. This raises a crucial question: is it possible to develop a truly time-optimal algorithm, specifically an $O(k)$-time algorithm, even if it requires much more space? In this section, we affirmatively answer this question for the rooted setting. We present an $O(k)$-time algorithm that utilizes $O(\Delta+\log k)$ bits of space per agent. However, the feasibility of an $O(k)$-time algorithm in the general setting remains open.
@@
We refer to the new algorithm as $\mathbf{RootedOpt}$ in this section. In $\mathbf{RootedDisp}$, with $O(\log(k+\Delta))$ bits of space, each settler $\psi(w)$ cannot memorize the exact set of settled neighbors of $w$. Instead, it only remembers the maximum $i$ such that the first $i$ neighbors of $w$ are settled. In contrast, $\mathbf{RootedOpt}$ allows each settler $\psi(w)$ to remember all settled neighbors of $w$ using $O(\Delta)$ bits, which significantly helps to eliminate an $O(\log\tau)$ factor from the time complexity. However, somewhat surprisingly, both the design of the new algorithm and the analysis of its execution time are non-trivial.
@@
Below, we outline the modifications made to $\mathbf{RootedDisp}$ to obtain $\mathbf{RootedOpt}$. We expect that most readers will grasp the behavior of $\mathbf{RootedOpt}$ simply by reviewing the following key differences, while we provide the pseudocode in Algorithm 6.
@@
  • In $\mathbf{RootedOpt}$, each settler $\psi(w)$ maintains an array variable $\psi(w).\mathtt{checked}$ of size $\delta_{w}$. Each element of $\psi(w).\mathtt{checked}[i]$ takes a value from the set $\{0,1,\bot\}$. The assignment $\psi(w).\mathtt{checked}[i]=0$ (respectively, $1$) indicates that the neighbor $N(w,i)$ is unsettled (respectively, settled). The value $\bot$ is utilized exclusively during the probing process, meaning that the neighbor $N(w,i)$ has yet to be checked for its settled status. For any $c\in\{0,1,\bot\}$, we define $U_{c}(u)=\{N(u,i)\mid i\in[0,\delta_{w}-1],\psi(w).\mathtt{checked}[i]=c\}$.
@@
  • Consider that the unique leader $a_{\mathrm{max}}$ invokes the probing process $\mathtt{Probe}()$ at a node $w$. (Remember that $a_{\mathrm{max}}$ is the unique leader that has the maximum identifier at the beginning of the execution.) In $\mathbf{RootedDisp}$, the probing process terminates as soon as any agent finds an unsettled neighbor. However, in $\mathbf{RootedOpt}$, the process only ends when all of $w$’s neighbors are probed or when at least $\ell$ unsettled neighbors are found, where $\ell$ is the number of explorers. In the former case, the probing process is now complete: $U_{\bot}(w)$ is empty, and $U_{0}(w)$ equals the set of unsettled neighbors of $w$. In the latter case, the explorers go to distinct $\ell$ unsettled nodes and settle there, thereby achieving dispersion.
@@
  • Consider an agent $a$ making a round trip $w\to u\to w$ during $\mathtt{Probe}()$, where $u=N(w,p)$ for some $p\in[0,\delta_{w}-1]$. If $a$ does not encounter a settler at $u$, it simply sets $\psi(w).\mathtt{checked}[i]$ to 00. On the other hand, if a settler is found at $u$, $a$ sets $\psi(w).\mathtt{checked}[i]$ to $1$ and additionally sets $\psi(u).\mathtt{checked}[q]$ to $1$, where $q\in[0,\delta_{u}-1]$ is the port number such that $w=N(u,q)$. This modification ensures that $U_{0}(u)$ remains equal to the set of unsettled neighbors of $u$ when the explorers go back to $u$.
@@
  • In $\mathbf{RootedDisp}$, the leader $a_{\mathrm{max}}$ invokes $\mathtt{Probe}()$ after each forward or backward move. However, in $\mathbf{RootedOpt}$, $a_{\mathrm{max}}$ only invokes $\mathtt{Probe}()$ after making a forward move. This change does not compromise the correctness of $\mathbf{RootedOpt}$ because when $a_{\mathrm{max}}$ makes a backward move to a node $w$, $\psi(w)$ accurately remembers its unsettled neighbors due to the modification mentioned earlier.
@@
One might think that the probing process $\mathtt{Probe}()$ in $\mathbf{RootedOpt}$ could take longer time than in $\mathbf{RootedDisp}$, as it only finishes after all neighbors of the current location have been probed or after finding $\ell$ unsettled neighbors, where $\ell$ is the number of explorers. Particularly, there seems to be a concern that during the probing at a node $w$, the number of agents, excluding $\psi(w)$, may not always double: this event occurs when some agents discover an unsettled neighbor. Despite that, we deny this conjecture at least asymptotically, that is, we have the following lemma.
@@
Assume that $a_{\mathrm{max}}$ invokes $\mathtt{Probe}()$ at node $w$ during the execution of $\mathbf{RootedOpt}$, and exactly $\ell$ explorers including $a_{\mathrm{max}}$ exists at the time. Then, $\mathtt{Probe}()$ finishes within $O(1)+\max(0,2\lceil\log\tau-\log\ell\rceil)$ time.
@@
If $\ell\geq\tau$, we have $\ell\geq\min(\Delta,k)=\Delta$ because $\ell<k$. Then, the lemma trivially holds: $\mathtt{Probe}()$ finishes in a constant time. Thus, we consider the case $\ell<\tau$. Let $t$ be the time step at which $a_{\mathrm{max}}$ invokes $\mathtt{Probe}()$ at a node $w$, and let $z=\lceil\log\tau-\log\ell\rceil+1$. It suffices to show that $U_{\bot}(w,t^{\prime})=\emptyset$ or $|U_{0}(w,t^{\prime})|\geq\ell$ holds for some $t^{\prime}\in[t,t+2z+O(1)]$. Assume for contradiction that this does not hold. For any $r\in[0,z]$, we define $f(r)=X_{r}\cdot 2^{r-1}+Y_{r}$, where $X_{r}=|U_{0}(w,t+2r)|$ and $Y_{r}=|A(w,t+2r)\setminus\{\psi(w)\}|$. By definition, $f(0)=0\cdot 2^{0-1}+\ell=\ell$. Under the above assumption, for any $r=0,1,\dots,z-1$, the agents in $A(w,t+2r)$ move to distinct neighbors in $U_{\bot}(w,t+2r)$ in time step $t+2r$, and bring back all settlers they find, at most one for each neighbor, in time step $t+2r+1$. Let $\alpha$ be the number of those settlers i.e.$\alpha=Y_{r+1}-Y_{r}$. Note that $X_{r+1}-X_{r}=Y_{r}-\alpha$ holds here. Then, irrespective of $\alpha$, we obtain
$f(r+1)$ $=X_{r+1}\cdot 2^{r}+Y_{r+1}=(X_{r}+Y_{r}-\alpha)\cdot 2^{r}+Y_{r}+\alpha$
$=X_{r}\cdot 2^{r}+(2^{r}-1)(Y_{r}-\alpha)+2\cdot Y_{r}\geq 2(X_{r}\cdot 2^{r-1}+Y_{r})=2f(r),$
where we use $2^{r}\geq 1$ and $Y_{r}-\alpha=X_{r+1}-X_{r}\geq 0$ in the above inequality. Therefore, we have $f(z)\geq\ell\cdot 2^{z}$, whereas we have assumed (for contradiction) that $X_{z}=|U_{0}(w,t+2z)|<\ell$, thus $f(z)=X_{z}\cdot 2^{z-1}+Y_{z}\leq(\ell-1)2^{z-1}+Y_{z}$ holds. This yields $|A(w,t+2z)\setminus\{\psi(w)\}|=Y_{z}\geq\ell\cdot 2^{z}-(\ell-1)2^{z-1}=(\ell+1)2^{z-1}\geq(\ell+1)\tau/\ell>\tau$. Since $\tau=\min(\Delta,k)$, we have $\tau=\Delta$ or $\tau=k$. In the former case, $Y_{k}>\Delta$ agents at $w$ are enough to visit all neighbors in $U_{\bot}(w,t+2z)$ in time step $t+2z$, thus $U_{\bot}(w,t+2z+2)=\emptyset$ holds, a contradiction. In the latter case, there are $|A(w,t+2z)|\geq k+2$ agents in $w$ at time step $t+2z$, a contradiction. Therefore, $U_{\bot}(w,t^{\prime})=\emptyset$ or $|U_{0}(w,t^{\prime})|\geq\ell$ holds at time step $t^{\prime}\leq t+2z+2$. ∎
@@
In the rooted setting, algorithm $\mathbf{RootedOpt}$ solves the dispersion problem within $O(k)$ time using $O(\Delta+\log k)$ bits of space per agent.